OpenStudy (anonymous):

The 18th term of a sequence is 243, and the 90th term is 516.6. What is the 125th term?

OpenStudy (anonymous):

The multiple choice answers given are: A. 624.6 B. 649.6 C. 653.4 D. 4,928.4

OpenStudy (rock_mit182):

is a geometric sequence ?

OpenStudy (anonymous):

It doesnt say... :/

OpenStudy (kc_kennylau):

common difference = (516.6-243)/(90-18) = 3.8

OpenStudy (kc_kennylau):

125th term = 516.6 + 3.8 * (125-90) = 649.6

OpenStudy (anonymous):

Thank you so much! I understand now

OpenStudy (kc_kennylau):

no problem :)